OpenStudy (anonymous):

. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 14x + 18y = -54 -4x – 14y = 42

OpenStudy (anonymous):

Multiply second one by 14/4 = 7/2 and add to get y.

jhonyy9 (jhonyy9):

14x +18y =-54 divide both sides by 2 and get 7x+9y=-27 -4x-14y=42 divide both sides by 2 and get -2x-7y=21 7x+9y=-27 multiply by 2 -2x-7y=21 multiply by 7 14x+18y=-54 -14x-49y=147 ---------------- assum 0 -31y=93 y=93/(-31) =-3 y=-3 -2x-7(-3)=21 -2x+21=21 -2x=0 x=0 y=-3

OpenStudy (anonymous):

good job kido